Jessica bought an antique candle holder for $300. The antique appreciates in value by 5.5% per year. When will the value of the candle holder be $3000? Round the answer to the nearest tenth.
Answer by stanbon(75887)

Jessica bought an antique candle holder for $300. The antique appreciates in value by 5.5% per year. When will the value of the candle holder be $3000? Round the answer to the nearest tenth.
-----
A(t) = P*(1.055)^t
----------
3000 = 300*1.055^t
1.055^t = 10
t = log(10)/log(1.055)
t = 1/0.0233
t = 43.0 years
--------------
Cheers,
Stan H.
------------